OpenStudy (smalls):

Choose the answer that correctly expresses this rational decimal in fraction form. 2.6

OpenStudy (anonymous):

2.6 can be converted into a fraction by placing it over 1: 2.6=2.6/1 Then I'm guessing you'd like to turn it into a nicer fraction with all whole numbers, so multiply the fraction by a number that makes 2.6 not a decimal: (2.6/1) * 5 = 13/5

OpenStudy (smalls):

thank yuh i was really stuckk

OpenStudy (anonymous):

No problem!